# Dark-Matter-Deficient Galaxies from Collisions: A New Probe of Bursty Feedback and Dark Matter Physics

### v2

@@ Abstract
High-velocity collisions between gas-rich ultra-diffuse galaxies present a promising formation channel for dark-matter-deficient galaxies (DMDGs). Using hydrodynamical simulations, we show that the progenitors’ baryonic binding energy, $|E_{\rm bind}|$, critically controls the outcome. Repeated potential fluctuations, e.g., from bursty feedback, inject energy and reduce $|E_{\rm bind}|$ by $\approx 15\%$, yielding fewer but substantially more massive DMDGs. By contrast, elastic self-interacting dark matter (SIDM) produces comparable cores without lowering $|E_{\rm bind}|$, resulting in negligible effect. This differs from the other DMDG formation scenario, where SIDM-induced cores enhance dark matter tidal stripping while keeping baryons compact and resilient to tidal effects. The contrasting roles of SIDM help distinguish feedback-formed halo cores from those created by SIDM. Among 15 paired simulation runs, 13 show higher DMDG masses in the weakened-binding case, and about two thirds exhibit $100\%$ mass enhancements. The simulations also predict systematically lower gas fractions due to sustained post-collision star formation, yielding a clean observational signature. Upcoming wide-field imaging (CSST, LSST), HI surveys (FAST), and kinematic follow-up will be crucial to test this scenario.
@@ 1 Introduction
In this work, we identify the baryonic binding energy of the progenitors, $|E_{\rm bind}|$, as the relevant control parameter, directly linking energy injection in dwarf halos to DMDG yields. We employ controlled hydrodynamical simulations of gas-rich ultra-diffuse galaxy (UDG) collisions to assess the potential of these systems as probes of energy injection in dwarf halos, positioning DMDG observations a new observational window for testing the underlying physics. Intriguingly, both baryonic feedback (J. F. Navarro et al., 1996; S. Mashchenko et al., 2006; F. Governato et al., 2010) and self-interacting dark matter (SIDM) (D. N. Spergel & P. J. Steinhardt, 2000; S. Tulin & H.-B. Yu, 2018)have been proposed as mechanisms for generating cores in halos, but their consequences diverge remarkably in collisional events. Future DMDG studies can thus inform both galaxy formation and dark sector physics.
2 Weakening gravitational binding by energy injection.
@@
Figure 1 shows the evolution of the density profile (left), gravitational potential (middle), and $|E_{\rm bind}|$ (right) in 2 Gyr. We compute $E_{\rm bind}$ by integrating the gas density over the total gravitational potential ($\Psi_{\rm tot}$), $E_{\rm bind}=4\pi\int drr^{2}\Psi_{\rm tot}\rho_{\rm gas}$. The results illustrate that impulsive feedback heats the inner orbits, leading to an expansion of centrally cored density profile, a shallowing of the central potential, and a reduction in the central baryonic binding energy. By $2$ Gyr, $|E_{\rm bind}|$ has decreased by over 10%. We will use collision simulations to show that even this modest reduction can substantially enhance the DMDG formation.
Table 1: Representative simulation benchmarks.
BM Profile $v_{r}$ (km/s) $|\frac{\Delta E_{\rm bind}}{E_{\rm bind}(\gamma=1)}|$ $\frac{M_{\rm b}(\gamma=0.1)}{M_{\rm b}(\gamma=1)}$
1a $\gamma=1$ 400 0.13 7.20
1b $\gamma=0.1$ 400
1c SIDM 400 0.08 1.83
2a $\gamma=1$ 400 0.15 5.74
2b $\gamma=0.1$ 400
3a $\gamma=1$ 280 0.15 2.53
3b $\gamma=0.1$ 280
Note. — The full simulation parameters and numerical results are provided in the Materials and Methods section.
Table 1 summarizes the configurations and results of representative simulations. The labels $\gamma=0.1$ and $\gamma=1$ correspond to simulations with and without weakened baryonic gravitational binding. We model the density profiles using an extended NFW profile with an inner slope controlled by the parameter $\gamma$ (see Materials and Methods for details).
3 Simulation setup
@@
where the central density slope is governed by $\gamma$, with smaller values of $\gamma$ corresponding to less gravitationally bound halos.
@@
We construct two contrasting sets of initial conditions, one with $\gamma=1$ (cuspy) and the other with $\gamma=0.1$ (cored, with weaker binding), while keeping the total halo mass and concentration fixed. The simulated halos span masses of $1-2\times 10^{10}~\rm M_{\odot}$ and concentrations in the range 4144-14. Assuming these halos host gas-rich UDGs, we model the baryonic content with a gas-only disk, as described previously, and vary the disk height between $0.15$ and $0.35~\rm kpc$. This maximizes the effectiveness of DMDG formation, as stars behave as a collisionless component during the collision and cannot be efficiently expelled from the progenitors. The gas-to-halo mass ratio is set to $0.1$ or $0.22$, with gas scale radii $r_{g}$ in the range 1.53.51.5-3.5 kpc. The pericentric distance is assumed to be $2\,\rm kpc$. These choices are motivated by observations of gas-rich UDGs, where internal turbulence and stellar feedback maintain gas densities below the threshold for molecular hydrogen formation, $n_{g}\lesssim 1~\mathrm{cm}^{-3}\approx 3\times 10^{7}~M_{\odot}\mathrm{kpc}^{-3}$, thereby suppressing star formation (A. Di Cintio et al., 2017; T. K. Chan et al., 2018). The population and properties of gas-rich UDGs have attracted considerable attention and been extensively explored in recent studies (E. Papastergis et al., 2017; F. Jiang et al., 2019; T. K. Chan et al., 2018; M. G. Jones et al., 2018; P. E. Mancera Piña et al., 2020; D. Kong et al., 2022).
@@
Equilibrium initial conditions are generated using DICE (V. Perret et al., 2014), with the gas disk stabilized by enforcing a Toomre $Q_{\rm gas}>1.5$ (A. Toomre, 1964; J. Binney & S. Tremaine, 2008; D. Williamson et al., 2016). The gas is initialized at $T=10^{4}$ K and metallicity $Z=0.1Z_{\odot}$. We evolve the collision simulations using Gadget-4 (V. Springel et al., 2021), incorporating radiative cooling of primordial gas and star formation based on the multiphase ISM model of V. Springel & L. Hernquist (2003). Stellar feedback is modeled as an effective pressure that prevents the gas from collapsing too quickly. To obtain the SIDM prediction, we set up a hydrodynamical CDM simulation using progenitors taken from the isolated SIDM simulation (Gadget-2) snapshot at 2 Gyr. We verified that switching off SIDM after this initialization yields a stationary profile in isolation. Further details are provided in the Supplementary Material.
@@
Our collision simulations build upon findings in previous studies, where dependencies of simulation outcomes on the collision velocities, progenitor halo mass and concentration, gas distribution, and numerical resolution were systematically explored. We adopt simulation setups similar to those in E.-j. Shin et al. (2020) and J. Lee et al. (2024), with relative collision velocities ($v_{r}$) ranging from $300-600~\rm km/s$. Our primary choice of $v_{r}=400~\rm km/s$ lies within the preferred range reported in the literature. We also perform simulations with $v_{r}=280$, $450$ and $500~\rm km/s$. The two colliding progenitors are initialized with identical conditions, placed at a separation of $60$ kpc along their relative velocity vector and with an impact parameter of $2$ kpc. We focus on contrasting paired samples with $\gamma=1$ and $\gamma=0.1$, testing them under a range of conditions. Table 2 summarizes the initial conditions and parameter setup in our simulations.
@@
Figure 2: Formation of DMDGs from the collision of gas-rich dwarf galaxies. The initial setup (left) has two progenitors approaching at a close separation. They collide within 0.2 Gyr (middle), displacing gas from the halo centers and triggering efficient star formation. By 5 Gyr (right), the expelled baryons have collapsed into several DMDGs that are far from their progenitors. The surface density distributions of dark matter (top), gas (middle), and stars (bottom) are presented at the corresponding snapshots for the 2a ($\gamma=1$) benchmark.
4 Collisional formation of DMDGs.
@@
Figure 2 shows simulation snapshots of the surface densities of dark matter (top), gas (middle), and stars (bottom) for the benchmark 2a ($\gamma=1$) in Table 1. Four stages of the collision are displayed. The first snapshot depicts the initial setup, with two progenitors approaching each other at close separation so that the simulation outcome primarily concerns the formation of DMDGs rather than the internal dynamics of the progenitors. The systems collide within 0.2 Gyr, displacing their gas from the halo centers and triggering efficient star formation. By 5 Gyr, the expelled baryons have already collapsed into several DMDGs and moved far from their progenitors.
With $v_{r}=400~\rm km/s$, most of the progenitor gas is expelled. In the cuspy case ($\gamma=1$), the steeper dark-matter potential exerts stronger tidal “winds,” causing the expelled gas to fragment into several small, low-mass condensations. In contrast, in the weakened-binding case ($\gamma=0.1$), the debris remains in contact more easily, allowing it to coalesce into fewer but more massive DMDGs.
To quantify how the tide acts on an extended gas debris, we evaluate on the mid plane the largest eigenvalue ($\lambda_{\rm max}$) of the tidal tensor $T\equiv-\nabla\nabla\Phi$, for which negative eigenvalues imply convergent relative acceleration. Figure 3 shows the head-on collisions between two identical progenitor halos, using parameters from the first benchmark in Table 1, at varying separations: $d=0$ (left), $4$ (middle), and $8~\rm kpc$ (right) for the $\gamma=1$ (top), $\gamma=0.1$ (middle), and SIDM (bottom) cases. We highlight the $\lambda_{\rm max}=0$ contour, within which all eigenvalues are negative and the tide is compressive.
@@
The $\gamma=0.1$ model develops broad, contiguous compressive regions around the origin across all separations, consistent with debris that readily coalesces into fewer, more massive clumps. A shallower inner slope renders the central potential more nearly harmonic, weakens shear, and suppresses stretching along the collision axis. By contrast, the cuspy $\gamma=1$ case exhibits only small, fragmented compressive islands around shear-dominated zones, consistent with the formation of multiple low-mass condensations. The SIDM core is encircled by a ring of stretching tides that inhibits gas confinement. The central compressive region still facilitates gas condensation, but the formation of DMDG is significantly suppressed compared with the $\gamma=0.1$ scenario.
@@
In a weakened gravitational binding scenario, gas is expected to escape more readily during collisions. To illustrate this, we reduce the collision velocity of the 3a and 3b benchmarks to $280~\mathrm{km/s}$ (see Table 1 for setup and results). The escaping gas mass reaches $3.8\times 10^{9}~M_{\odot}$ in the $\gamma=0.1$ case, compared to $3.3\times 10^{9}~M_{\odot}$ for $\gamma=1$. Both the stellar and gas masses bound to the resulting DMDGs are substantially higher in the $\gamma=0.1$ scenario, with enhancements of $240\%$ and $34\%$, respectively. The most massive DMDG formed in our simulations has a stellar mass of $1.68\times 10^{9}~M_{\odot}$, obtained in the $\gamma=0.1$, $v_{r}=280~\mathrm{km/s}$ run. These results demonstrate that reduced baryonic binding favors the formation of more massive DMDGs.
@@
Figure 4: Characteristics of the most massive DMDGs in 15 paired simulations. The bar chart summarizes results from contrasting simulations with inner slopes $\gamma=0.1$ and $\gamma=1$. Bars are arranged in ascending order of the DMDG mass ratio, $M_{b,\gamma=0.1}/M_{b,\gamma=1}$. The bar width scales with the relative change in binding energy, $(E_{{\rm bind},\gamma=1}-E_{{\rm bind},\gamma=0.1})/E_{{\rm bind},\gamma=1}$, while the color indicates the gas fraction, $f=M_{\rm gas}/(M_{\rm gas}+M_{\rm stars})$. In contrast, the increase in the DMDG masses is predominantly positive. In 9 out of the 15 simulations, this increase exceeds 100%, as reflected by the bars that rise above the line representing $\Delta M_{b}/M_{b,\gamma=1}=1$.
@@
Aside from the examples shown, we simulated 15 pairs of dwarf collisions with varied setups. Figure 4 presents the relative DMDG mass enhancements in the $\gamma=0.1$ and $\gamma=1$ cases, with bar widths proportional to the reduction in baryonic binding energy and colors indicating the gas fraction, $f=M_{\rm gas}/(M_{\rm gas}+M_{\rm stars})$. We find that 13 of the 15 simulations yield an increase in DMDG mass, with 9 of these cases showing an enhancement greater than 100%. The two exceptions occur for halos with extremely low concentrations ($c=4$) and $v_{r}=450~\mathrm{km/s}$, conditions that reduce the contrast between the two scenarios. Interestingly, the gas fractions in the $\gamma=0.1$ runs are systematically lower than in their $\gamma=1$ counterparts, demonstrating more sustained star formation following the collision. In this figure, we illustrate results considering only the most massive DMDGs. Our findings remain largely unchanged when considering the sum of multiple DMDG masses. The Materials and Methods section presents analogous results that include the total DMDG mass as well as the case with the gas cooling rate reduced by half. Additional details on post-collision star formation are also provided.
5 Discussion and Conclusion

### v3

@@ Abstract
High-velocity collisions between gas-rich ultra-diffuse galaxies present a promising formation channel for dark-matter-deficient galaxies (DMDGs). Using hydrodynamical simulations, we show that the progenitors’ baryonic binding energy, $|E_{\rm bind}|$, critically controls the outcome. Repeated potential fluctuations, e.g., from bursty feedback, inject energy and reduce $|E_{\rm bind}|$ by $\approx 15\%$, yielding fewer but substantially more massive DMDGs. By contrast, elastic self-interacting dark matter (SIDM) produces comparable cores without lowering $|E_{\rm bind}|$, perturbing DMDG masses without clear enhancement. This differs from what happens in host halos, where SIDM-induced cores enhance dark matter tidal stripping while keeping baryons compact and resilient to tidal effects. The contrasting roles of SIDM may provide a means to distinguish feedback-formed halo cores from those created by SIDM. Among 15 paired simulation runs, 13 show higher DMDG masses in the weakened-binding case, and about two thirds exhibit $100\%$ mass enhancements. The simulations also predict systematically lower gas fractions due to sustained post-collision star formation, yielding a clean observational signature. Upcoming wide-field imaging (CSST, LSST), HI surveys (FAST), and kinematic follow-up will be crucial to test this scenario.
@@ I Introduction
In this work, we identify the baryonic binding energy of the progenitors, $|E_{\rm bind}|$, as the relevant control parameter, directly linking energy injection in dwarf halos to DMDG yields. We employ controlled hydrodynamical simulations of gas-rich ultra-diffuse galaxy (UDG) collisions to assess the potential of these systems as probes of energy injection in dwarf halos, positioning DMDG observations a new observational window for testing the underlying physics. As both baryonic feedback (J. F. Navarro et al., 1996; S. Mashchenko et al., 2006; F. Governato et al., 2010; J. D. Burger et al., 2022) and self-interacting dark matter (SIDM) (D. N. Spergel & P. J. Steinhardt, 2000; S. Tulin & H.-B. Yu, 2018) have been proposed as mechanisms for generating cores in halos, we also set up simulations to test the capabilities of feedback and SIDM in boosting DMDG formations. Our results reveal distinct differences in their effects on DMDG formation. As a result, these two core formation mechanisms are expected to produce different DMDG populations in isolated and host environments. Measuring these populations offers an observational method to advance understanding of galaxy formation and dark sector physics.
II Weakening gravitational binding by energy injection.
@@
where the central density slope is governed by $\gamma$, with smaller values of $\gamma$ corresponding to less gravitationally bound halos. Figure 1 shows the evolution of the density profile (left), gravitational potential (middle), and $|E_{\rm bind}|$ (right) in 2 Gyr. We compute $E_{\rm bind}$ by integrating the gas density over the total gravitational potential ($\Psi_{\rm tot}$), $E_{\rm bind}=4\pi\int drr^{2}\Psi_{\rm tot}\rho_{\rm gas}$. The results illustrate that impulsive feedback heats the inner orbits, leading to an expansion of centrally cored density profile, a shallowing of the central potential, and a reduction in the central baryonic binding energy. By $2$ Gyr, $|E_{\rm bind}|$ has decreased by over 10%. We will use collision simulations to show that even this modest reduction can substantially enhance the DMDG formation.
Table 1: Galaxy collision simulations with Gadget-4: initial collision configurations and properties of the resulting dark-matter-deficient galaxies (DMDGs).
BM $M_{200,\,\rm DM}$ $M_{\rm gas}$ $v_{r}$ $R_{s}$ $c$ $\gamma$ $|E_{\rm bind}|$ $n$ $M_{\rm star}$ $M_{\rm gas}$
($10^{10}\,M_{\odot}$) ($10^{10}\,M_{\odot}$) (km s-1) (kpc) $E_{0}$ ($10^{8}\,M_{\odot}$) ($10^{8}\,M_{\odot}$)
1a 1.5 0.15 $400$ 22 $14$ 1 8.75 1 0.38 1.49
1b 0.1 7.58 1 8.61 4.85
1c SIDM 9.47 1 0.46 2.97
2a 2 0.2 $400$ 22 $14$ 1 14.73 4 0.10 2.13
2b 0.1 12.49 1 6.09 6.72
2c SIDM 15.92 2 0.69 3.70
3a 2 0.2 $280$ 22 $14$ 1 14.73 2 4.89 3.70
3b 0.1 12.49 1 16.80 4.97
3c SIDM 15.92 3 2.68 2.98
4a 1.0 0.1 $400$ $1.5$ 44 1 3.53 1 0.15 4.13
4b 0.1 2.87 1 0.30 5.54
4c SIDM 4.06 1 0.04 1.39
5a 1.5 0.15 $400$ 22 44 1 7.17 1 6.96 8.32
5b 0.1 5.82 1 9.61 5.05
5c SIDM 8.28 1 3.74 4.24
6a 1.5 0.22 $400$ 22 44 1 10.99 2 16.84 6.02
6b 0.1 9.28 1 20.70 7.03
6c SIDM 12.64 2 4.08 3.79
7a 1.5 0.22 $450$ 22 44 1 10.99 2 2.39 7.74
7b 0.1 9.28 2 3.23 3.96
7c SIDM 12.64 1 0.18 0.68
8a 1.5 0.15 $400$ 22 77 1 7.87 3 1.62 3.38
8b 0.1 6.47 1 11.15 3.91
8c SIDM 8.74 3 0.25 3.33
9a 1.5 0.22 $400$ 22 77 1 11.93 3 4.66 4.42
9b 0.1 10.21 2 18.33 5.57
9c SIDM 13.18 1 15.79 3.60
10a 1.5 0.22 $280$ 22 77 1 11.93 2 12.71 3.15
10b 0.1 10.21 1 25.90 7.27
10c SIDM 13.18 1 20.81 6.36
11a 1.5 0.15 $280$ 22 $14$ 1 8.75 1 0.22 2.14
11b 0.1 7.58 1 8.38 5.20
11c SIDM 9.47 3 3.50 3.14
12a 1.5 0.22 $400$ 22 $14$ 1 13.56 3 3.12 5.88
12b 0.1 11.68 1 13.15 5.23
12c SIDM 14.53 2 4.90 5.32
13a 2 0.2 $400$ 22 44 1 11.53 1 14.64 6.60
13b 0.1 9.75 1 19.49 5.88
13c SIDM 13.20 2 7.60 4.26
14a 2 0.2 $400$ 22 77 1 12.81 2 0.78 3.01
14b 0.1 10.83 1 13.97 5.53
14c SIDM 14.18 1 11.60 5.19
15a 5 0.5 $500$ $3.5$ $14$ 1 78.10 3 13.81 4.98
15b 0.1 66.40 1 26.08 5.11
15c SIDM 84.24 4 11.38 4.89
Table 1 summarizes the configurations and results of representative simulations. The labels $\gamma=0.1$ and $\gamma=1$ correspond to simulations with and without weakened baryonic gravitational binding. We model the density profiles using an extended NFW profile with an inner slope controlled by the parameter $\gamma$.
III Collision Simulation Setup
@@
As described above, we construct two contrasting sets of initial conditions, one with $\gamma=1$ (cuspy) and the other with $\gamma=0.1$ (cored, with weaker binding), while keeping the total halo mass and concentration fixed. The detailed simulation parameters are summarized in Table 1. We adopt a particle mass of $55\times 10^{4}\,\rm M_{\odot} 11$1For the final sample with $M_{200,\rm DM}=5\times 10^{10}\,M_{\odot}$, the particle mass is $3.95\times 10^{4}\,\rm M_{\odot}$. and a gravitational softening length of $\epsilon=0.2\rm\,kpc$ that satisfies $\epsilon\approx\frac{4r_{200}}{\sqrt{N_{200}}}$ (C. Power et al., 2003). The simulated halos span masses of $1-2\times 10^{10}~\rm M_{\odot}$ and concentrations in the range 4144-14. Assuming these halos host gas-rich UDGs, we model the baryonic content with a gas-only disk, as described previously, and vary the disk height between $0.15$ and $0.35~\rm kpc$. This maximizes the effectiveness of DMDG formation, as stars behave as a collisionless component during the collision and cannot be efficiently expelled from the progenitors. The gas-to-halo mass ratio is set to $0.1$ or $0.22$, with gas scale radii $r_{g}$ in the range 1.53.51.5-3.5 kpc. These choices are motivated by observations of gas-rich UDGs, where internal turbulence and stellar feedback maintain gas densities below the threshold for molecular hydrogen formation, $n_{g}\lesssim 1~\mathrm{cm}^{-3}\approx 3\times 10^{7}~M_{\odot}\mathrm{kpc}^{-3}$, thereby suppressing star formation (A. Di Cintio et al., 2017; T. K. Chan et al., 2018). The population and properties of gas-rich UDGs have attracted considerable attention and been extensively explored in recent studies (E. Papastergis et al., 2017; F. Jiang et al., 2019; T. K. Chan et al., 2018; M. G. Jones et al., 2018; P. E. Mancera Piña et al., 2020; D. Kong et al., 2022).
@@
Equilibrium initial conditions are generated using DICE (V. Perret et al., 2014), with the gas disk stabilized by enforcing a Toomre $Q_{\rm gas}>1.5$ (A. Toomre, 1964; J. Binney & S. Tremaine, 2008; D. Williamson et al., 2016). The gas is initialized at $T=10^{4}$ K and metallicity $Z=0.1Z_{\odot}$. We evolve the collision simulations using Gadget-4 (V. Springel et al., 2021), incorporating radiative cooling of primordial gas and star formation based on the multiphase ISM model of V. Springel & L. Hernquist (2003). The averaged thermodynamic impact of stellar feedback is modeled as an effective pressure that prevents the gas from collapsing too quickly. To obtain the SIDM prediction, we set up a hydrodynamical CDM simulation using progenitors taken from the isolated SIDM simulation (Gadget-2) snapshot at 2 Gyr. We verified that switching off SIDM after this initialization yields a stationary profile in isolation. Further details are provided in Appendix A.
@@
Our collision simulations build upon findings in previous studies, where dependencies of simulation outcomes on the collision velocities, progenitor halo mass and concentration, gas distribution, and numerical resolution were systematically explored. We adopt simulation setups similar to those in E.-j. Shin et al. (2020) and J. Lee et al. (2024), with relative collision velocities ($v_{r}$) ranging from $300-600~\rm km/s$. Our primary choice of $v_{r}=400~\rm km/s$ lies within the preferred range reported in the literature. We also perform simulations with $v_{r}=280$, $450$ and $500~\rm km/s$. The two colliding progenitors are initialized with identical conditions, placed at a separation of $60$ kpc along their relative velocity vector and with an impact parameter of $2$ kpc. We focus on contrasting paired samples with $\gamma=1$ and $\gamma=0.1$, as well as an SIDM scenario, testing them under a range of conditions. All the collision simulations in this work use the same module for gas and stellar physics. It follows that all the simulation outcomes differ only because of the initial conditions. The initial condition files for the simulations listed in Table 1 and the corresponding configuration files for Gadget-4 are publicly available at https://pan.cstcloud.cn/s/xbilapRkS5Q.
@@
Figure 2: Formation of DMDGs from the collision of gas-rich dwarf galaxies. The initial setup (left) has two progenitors approaching at a close separation. They collide within 0.2 Gyr (middle), displacing gas from the halo centers and triggering efficient star formation. By 5 Gyr (right), the expelled baryons have collapsed into several DMDGs that are far from their progenitors. The surface density distributions of dark matter (top), gas (middle), and stars (bottom) are presented at the corresponding snapshots for the 2b ($\gamma=1$) benchmark.
IV Collisional formation of DMDGs.
@@
Figure 2 shows simulation snapshots of the surface densities of dark matter (top), gas (middle), and stars (bottom) for the benchmark 2a ($\gamma=1$) in Table 1. Three stages of the collision are displayed. The first snapshot depicts the initial setup, with two progenitors approaching each other at close separation so that the simulation outcome primarily concerns the formation of DMDGs rather than the internal dynamics of the progenitors. The systems collide within 0.2 Gyr, displacing their gas from the halo centers and triggering efficient star formation. By 5 Gyr, the expelled baryons have already collapsed into several DMDGs and moved far from their progenitors. As shown in Table 1, progenitors with $\gamma=0.1$ systematically produce more massive DMDGs. With $v_{r}=400~\rm km/s$, most of the progenitor gas is expelled. To understand the origin of these differences, we quantify the role of tidal effect and star formation efficiency in regulating gas separation and collapse during the collision.
@@
To quantify how the tide acts on an extended gas debris, we evaluate on the mid plane the largest eigenvalue ($\lambda_{\rm max}$) of the tidal tensor $T\equiv-\nabla\nabla\Phi$, for which negative eigenvalues imply convergent relative acceleration. A positive $\lambda_{\rm max}$ corresponds to stretching tides, which promote the separation and fragmentation of expelled gas into multiple low-mass condensations. Figure 3 shows the head-on collisions between two identical progenitor halos, using parameters from the first benchmark in Table 1, at varying separations: $d=0$ (left), $4$ (middle), and $8~\rm kpc$ (right) for the $\gamma=1$ (top), $\gamma=0.1$ (middle), and SIDM (bottom) cases. We highlight the $\lambda_{\rm max}=0$ contour, within which all eigenvalues are negative and the tide is compressive. The $\gamma=0.1$ model develops broad, contiguous compressive regions around the origin across all separations, consistent with debris that readily coalesces into fewer, more massive clumps. A shallower inner slope renders the central potential more nearly harmonic, weakens shear, and suppresses stretching along the collision axis. By contrast, the cuspy $\gamma=1$ case exhibits only small, fragmented compressive islands around shear-dominated zones, consistent with the formation of multiple low-mass condensations. The SIDM core is encircled by a ring of stretching tides that inhibits gas confinement. The central compressive region still facilitates gas condensation, but the formation of DMDG is significantly suppressed compared with the $\gamma=0.1$ scenario.
As tidal forces are exerted at scales larger than the clumps to form, they operate as a “wind”, which shallows or steepens the gravitational potential as it sweeps over a region that would collapse, depending on whether the tide is repulsive or compressive. Consequently, in the cuspy case ($\gamma=1$), the stronger repulsive tidal wind causes the expelled gas to fragment into several small, low-mass condensations. In contrast, in the weakened-binding case ($\gamma=0.1$), the debris remains in contact more easily, allowing it to coalesce into fewer but more massive DMDGs. In Appendix B, we provide further evidence showing that the tidal wind is sufficiently strong to affect the gas condensation right after the collision.
@@
In a weakened gravitational binding scenario, gas is expected to escape more readily during collisions. To illustrate this, we reduce the collision velocity of the 3a and 3b benchmarks to $280~\mathrm{km/s}$ (see Table 1 for setup and results). The escaping gas mass reaches $3.8\times 10^{9}~M_{\odot}$ in the $\gamma=0.1$ case, compared to $3.3\times 10^{9}~M_{\odot}$ for $\gamma=1$. Both the stellar and gas masses bound to the resulting DMDGs are substantially higher in the $\gamma=0.1$ scenario, with enhancements of $240\%$ and $34\%$, respectively. The most massive DMDG formed in our simulations has a stellar mass of $2.60\times 10^{9}~M_{\odot}$ in 15b, obtained in the $\gamma=0.1$, $v_{r}=500~\mathrm{km/s}$ run. These results demonstrate that reduced baryonic binding favors the formation of more massive DMDGs.
@@
Figure 4: Characteristics of the most massive DMDGs in 15 paired simulations. The bar chart summarizes the ratios of the most massive DMDG masses in the $\gamma=0.1$ (left) and SIDM (right) scenarios, each normalized relative to the $\gamma=1$ case. Bars are arranged in ascending order of the DMDG mass ratio, $M_{b,\gamma=0.1}/M_{b,\gamma=1}$ or $M_{b,SIDM}/M_{b,\gamma=1}$. The bar width scales with the relative change in binding energy, $|\Delta E_{{\rm bind}}/E_{{\rm bind},\gamma=1}|$, while the color indicates the gas fraction, $f=M_{\rm gas}/(M_{\rm gas}+M_{\rm stars})$. With lowered gravitational binding, the increase in the DMDG masses is predominantly positive. In 9 out of the 15 simulations, this increase exceeds 100%, as reflected by the bars that rise above the line representing $\Delta M_{b}/M_{b,\gamma=1}=1$. In the SIDM scenario, mass increases and decreases occur with nearly equal frequency. There is also a weak anti-correlation between the magnitude of mass change and gas fraction.
@@
Aside from the examples shown, we simulated 15 sets of dwarf collisions with varied setups. Figure 4 presents the most massive DMDG mass in the $\gamma=0.1$ (left) or SIDM (right) cases, relative to the $\gamma=1$ case, with bar widths proportional to the reduction in baryonic binding energy and colors indicating the gas fraction, $f=M_{\rm gas}/(M_{\rm gas}+M_{\rm stars})$. In the $\gamma=0.1$ cases on the left panel, 13 out of the 15 simulations yield an increase in DMDG mass, with 9 of these cases showing an enhancement greater than 100%. The two exceptions occur for halos with extremely low concentrations ($c=4$) and $v_{r}=450~\mathrm{km/s}$, conditions that reduce the contrast between the two scenarios. On the right panel, we find that half of the most massive DMDGs in SIDM have lower masses than the $\gamma=1$ ones. The other half have increased masses with amplitudes close to those of the decreased ones, which are overall less significant than the increases in the $\gamma=0.1$ cases. These results agree with our expectations from the tidal wind analyses. Interestingly, the gas fractions in the $\gamma=0.1$ runs are systematically lower than in their $\gamma=1$ counterparts, demonstrating more sustained star formation following the collision in the $\gamma=0.1$ cases. In SIDM, the gas fraction depicts an anti-correlation with the gas mass changes $\Delta M_{b}$, where the cases with increased DMDG masses have a lowered gas fraction, while the decreased mass cases have an increased gas fraction. In combination with the tidal analysis in Figure 3, we find that the compressive (blue colors) and repulsive (red colors) tides may have caused the decrease and increase in the gas fractions. While both tides appear in SIDM, the compressive one dominates in the $\gamma=0.1$ case. Compressive tides can compress gas clumps, drive stronger star formation, and reduce gas fractions after a few Gyr. We provide details on the post-collisional star formation in Appendix C. In addition, we present in Appendix D analogous results that include the total DMDG mass and the case with the gas cooling rate reduced by half. Our findings remain largely unchanged in these alternative scenarios.
V DMDGs from tidal stripping in low-velocity collisions
So far in this work, we have focused on high-velocity collisions in which DMDGs form outside progenitor halos. Our results show that a lower $E_{\rm bind}$ efficiently promotes DMDG formation, whereas SIDM cannot. Intriguingly, the formation of DMDGs has been explored in the context of SIDM, but for satellite galaxies (D. Yang et al. (2020); Z.-C. Zhang et al. (2025a); Z. C. Zeng et al. (2025)). For example, D. Yang et al. (2020) shows that a DF2-like DMDG in the NGC1052 host halo can form more easily under SIDM, where the cored density profile boosts the tidal stripping of dark matter, making the satellite galaxy deficient in dark matter. Such analyses suggest that SIDM also increases DMDG production, but in a manner different from stellar feedback. Therefore, a population of DMDGs, both in isolation and in hosts, offers a novel window to differentiate between the two core formation mechanisms.
To explore how stellar feedback and SIDM-generated cores differ in producing DMDGs, here we follow the setup of D. Yang et al. (2020), placing our progenitor halo systems on orbits evolving in a host potential. The host galaxy consists of a dark matter halo and a stellar component. The halo follows an NFW density profile with $M_{200}=1.1\times 10^{13}~{\rm M_{\odot}}$ and $r_{s}=80~\rm kpc$, while the stellar component is modeled as a Hernquist profile (L. Hernquist, 1990) with scale density $\rho_{h}=1.1\times 10^{10}~\rm M_{\odot}\,kpc^{-3}$ and scale radius $r_{h}=1.2~\rm kpc$. We set up three simulations for the $\gamma=1$, $\gamma=0.1$, and SIDM scenarios, taking the 2a, 2b, and 2c benchmarks from Table 1 as initial satellites and put them at the same orbital apocenter, $380~\rm kpc$ from the host center, with a tangential velocity of $27~\rm km~s^{-1}$. The simulations use Gadget-2 and treat all baryon particles as collisionless. To accurately model the tidal evolution, we consider $\gtrsim 10^{6}$ dark matter simulation particles and set the masses of baryon particles to be equal. The SIDM simulation adopts $\sigma/m=5~\rm cm^{2}/g$, a value widely adopted in the literature to address small-scale challenges. The initial core from the 2c benchmark can be interpreted as coming from pre-infall gravothermal evolution.
Figure 5: Tidal evolution of satellite systems in a $10^{13}~\rm M_{\rm\odot}$ host halo, with the initial conditions taken from the 2a, 2b, and 2c benchmarks from Table 1 for the $\gamma=1$ (blue), $\gamma=0.1$ (red), and SIDM (orange) scenarios. Left: Evolution of the bound dark matter (solid) and stellar (dashed) masses. The region giving rise to a close DF2 analog is shaded in light purple. Right: Evolution of the stellar-to-halo mass ratio. DMDGs are often selected by $M_{\rm star}/M_{\rm DM}>1$. The shaded area corresponds to the same region as in the left panel.
Figure 5 presents the simulation results for the evolution of bound masses (left) and stellar-to-halo mass ratios (right). On the left panel, both the dark matter (solid) and the stellar (dashed) masses decrease along with the tidal evolution, with each steep mass loss corresponding to a pericentric passage. At the first of such passages, the $\gamma=0.1$ case shows already more significant mass losses in both dark matter and stars than the other two cases. This suggests that a lowered gravitational binding facilitates tidal stripping. The SIDM case is similar to the $\gamma=1$ case, since their potentials look similar beyond the core region. As the satellites’ tidal radii shrink along with their mass loss, tidal stripping removes masses from progressively inner regions, and we start to see that SIDM cores accelerate tidal stripping. On the right panel, the stellar-to-halo mass ratio increases more rapidly in $\gamma=0.1$ right after the first pericentric passage, exceeding one in about 5 Gyrs. The growth continues and passes 10 around 12 Gyr, when the SIDM case starts to show a surge in the ratio above 1.
Both the lowered gravitational binding and SIDM scenarios produce DMDGs within the age of the universe, but the required times and DMDG masses differ significantly. A DF2 analog of mass around $10^{8}~\rm M_{\odot}$ forms in bout 8 Gyr in the $\gamma=0.1$ case, and we shade the region of relevance in light purple. Such an analog appears in SIDM at around 13 Gyr, with a much lower stellar-to-halo mass ratio. These results reveal the different capabilities of stellar feedback and SIDM in producing DMDGs through tidal stripping.
VI Discussion and Conclusion